### v3

# AI Copilots for Reproducibility in Science: A Case Study

Abstract
Open science initiatives seek to make research outputs more transparent, accessible, and reusable, but ensuring that published findings can be independently reproduced remains a persistent challenge. This paper introduces OpenPub, an AI-powered platform that supports researchers, reviewers, and readers through a suite of modular copilots focused on key open science tasks. In this work, we present the Reproducibility Copilot, which analyzes manuscripts, code, and supplementary materials to generate structured Jupyter Notebooks and recommendations aimed at facilitating computational, or “rote”, reproducibility. We conducted feasibility tests using previously studied research papers with known reproducibility benchmarks. Results indicate that OpenPub can substantially reduce reproduction time - from over 30 hours to about 1 hour - while achieving high coverage of figures, tables, and results suitable for computational reproduction. The system systematically detects barriers to reproducibility, including missing hyperparameters, undocumented preprocessing steps, and incomplete or inaccessible datasets. While preliminary, these findings suggest that AI-driven tools can meaningfully reduce the burden of reproducibility efforts and contribute to more transparent and verifiable scientific communication. The modular copilot architecture also provides a foundation for extending AI assistance to additional open science objectives beyond reproducibility.
Introduction
The scientific enterprise is a highly collaborative effort, with research conducted by groups of scientists working in separate labs that are often scattered around the world. As such, scientific progress is critically dependent on the accurate, truthful communication and curation of technical results. The scientific publication process, in particular the journal publication process, has been the primary means of communicating technical results for hundreds of years.
The desire to share and reuse scientific data and software has grown in recent years as funders (e.g., National Science Foundation [1, 2, 3]) and publishers (e.g., American Geophysical Union [4]) have introduced open science policies that emphasize reproducibility, and as scientists increasingly recognize the benefits of open science [5, 6]. Consequently, the last decade has seen a proliferation of frameworks (e.g., electronic notebooks and containers) to promote open source resources, collaboration among scientists, and sharing of science products.
The advent of the Web also ushered in the Open Access Movement, lowering the barrier of subscription charges, and the Web environment has also encouraged experimentation with a wide variety of approaches to reviewing and feedback. More recently, the concept of “Open Science,” which broadly refers to the open and transparent sharing of data, physical samples, software, models, knowledge, etc., has also become a rallying point for the scientific community [7].
As discussed by Gil et al. (2016), the practice of open science can require additional time and work by researchers, such as documenting code and data so that analyses can be reproduced [8]. However, AI can now have a high impact in helping researchers in this endeavor. Indeed, the striking new advances in AI over the past few years, largely driven by the success of large language models (LLMs), have created new opportunities to improve how science is conducted and potentially revolutionize current paradigms for reviewing, curating, publishing, distributing, and sharing new technical results.
In this paper, we present a case study that shows how large the impact of AI can have on making science more open. To make that possible, we are developing OpenPub, an AI-enabled set of copilots for helping the different actors (authors, readers, reviewers, publishers, etc.) in the open science enterprise. In this study, we focus on a copilot that helps authors make their research more reproducible and that helps readers to reproduce the authors’ research. We developed an initial version of this copilot, and demonstrated the feasibility of the approach. For instance, we revisited a study on reproducibility by Gundersen et al. [9], and showed that for a paper that previously took 30 hours to reproduce, our copilot could produce recommendations which, if followed, would reduce the time to reproduce the paper to one hour.
Use Case: A Copilot for Reproducing Research
The “Reproducibility Copilot” that we developed serves two complementary functions: assisting authors in making their work more reproducible and guiding readers through the process of reproducing the work. The prototype was designed to operate on typical research artifacts (manuscripts, supplementary material, code, and datasets) and generate customized guidance, such as a Jupyter Notebook for readers or annotated comments in PDFs for authors.
In considering what problems the Reproducibility Copilot should tackle, we were influenced by the analysis of Gundersen et al. (2025), who analyzed the impact of different issues related to making science more open [9]. Their categorization included impactful issues, such as missing hyperparameter values for reproducing experiments, missing or undocumented code components, and incomplete data references. Using Gundersen et al. analysis as a guide, we designed a copilot capable of detecting and providing recommendations to fix a spectrum of problems.
Figure 1 provides a high-level overview of the architecture of the system, which is designed to support both the authors and the readers in a coordinated back-and-forth manner. This dual role design allows the system to anticipate the reader’s perspective, identify elements that can be difficult to interpret or reproduce, and generate a targeted output for the reader (e.g., a Jupyter Notebook to guide reproduction) while also producing actionable recommendations for the author (e.g., comments in the manuscript and code annotations).
Fig 1: System architecture of OpenPub, our AI-driven Reproducibility Copilot, illustrating its dual-path design that simultaneously supports authors and readers.
One of the first steps of the Reproducibility Copilot is to generate a reproducibility-focused Jupyter Notebook for the reader. This Notebook is structured around the logical flow of the original study, using placeholders where information is missing from the authors. These placeholders serve not only to guide the reader through incomplete parts of the research, but also to inform the author about where clarifications or additions are needed. Anchoring the LLM analysis on this Notebook helps focus the checks on the most essential elements for reproduction.
The generated Notebook, shown in Figure 2, begins with a clear structure to reproduce the figures and tables of the study. By organizing the reader’s process and clarifying where input is needed, the Notebook serves as a scaffold rather than a replacement for the author’s work. In Figure 3, we show an example cell that corresponds to a figure from Rodriguez and Laio (2014) [10]. The system has added a placeholder to indicate that the code to generate the plot is missing, and uses this signal to later generate a recommendation to the author to supply the relevant code.
Fig 2: Opening section of the Jupyter Notebook automatically generated by the system to guide readers in reproducing the experiments from Rodriguez and Laio (2014) [10].
Fig 3: Excerpt from the system-generated Notebook for reproducing Figure 4 from Rodriguez and Laio (2014) [10]. Since the original code for this figure was not provided, the system inserts minimal placeholder content to signal the missing implementation.
Once the Notebook is generated, the system applies a series of dedicated modules to identify and flag missing or ambiguous information. Through successive design iterations, four modules were designed and implemented in our prototype:
  1. 1.
    Hyperparameter Checker: Analyzes code and documentation to identify critical hyperparameter values that are missing or ambiguous. Recommends clarifications to the authors.
  2. 2.
    Dataset Checker: Scans the manuscript and code for data usage, verifying the presence of direct links to datasets. Flags missing datasets and proposes resolution steps.
  3. 3.
    Code Checker: Examines code to detect missing snippets essential for experiment replication.
  4. 4.
    Documentation Checker: Assesses inline comments and code structure to evaluate readability and comprehensibility. Generates clarifications for unclear segments.
Each of these modules is implemented using GPT-4o [11] and a 2-step prompt strategy. For the first step, a prompt was designed to list candidates in each specific module. For instance, in the case of the Hyperparameter Checker, this first step consisted in listing all the hyperparameters mentioned in the manuscript and code. The second step then uses a specific prompt to reduce the list according to the issue at hand and provides some recommendations to the authors. In the case of the Hyperparameter Checker, this consisted in keeping the hyperparameters for which some values are not clearly provided for some experiments. Given the non-deterministic nature of GPT-4o, these checks were performed five times, and another prompt was used with GPT-4o to consolidate these results by taking the union of the final lists produced.
Once all module outputs are consolidated, the final list produced by each module is used by the system to annotate the manuscript PDF with targeted highlights and margin comments, and inserts inline feedback within code files. As the number of comments can be high, depending on the reproducibility issues in the paper, the corresponding feedback comments were deliberately kept minimal to avoid overwhelming the authors.
Results
A crucial result of our case study is the demonstration of a significant time reduction in reproducing scientific research. To demonstrate the practical impact of the system, we selected a paper by Rodriguez and Laio (2014) [10], which had previously been analyzed in detail by Gundersen et al. (2025) [9]. This paper, which introduced a new machine learning method for clustering, was chosen specifically because it was one of the studies successfully reproduced by Gundersen et al., confirming that replication of the results was possible. Furthermore, the reported reproduction time of 33 hours by Gundersen et al. was below the 40-hour cutoff used in their study. Since this reproduction effort did not reach the cutoff limit, the 33-hour figure reflects the complete time it took to fully reproduce the study, making it a reliable benchmark against which we could measure our system’s efficiency.
Although it took a researcher 33 hours to reproduce the experiments in the Gundersen et al. study, using our copilot system, we found that users could complete the task in a relatively short time. Specifically, we asked two individuals (who were unfamiliar with the reproduced paper, its methodology, and its data) to independently play the role of readers and use our copilot to replicate the study. The first user, a researcher with a Ph.D. in data science, completed the task in about 1 hour. The second, a somewhat less experienced user (a college student in computer science with substantial AI experience), also achieved similar performance (i.e., about 1 hour to reproduce the experiments), suggesting that the system makes reproducibility more accessible, not just more efficient.
These tests were carried out under the assumption that the author is fully compliant with the recommendations generated by our tool, allowing us to evaluate the effectiveness of the system in an ideal usage scenario. What this means is that the copilot generated suggestions for how to modify the code and documentation, then we (members of our project team) played the role of the authors, modifying and/or adding code, data, and documentation according to the system’s suggestions. To emulate the role of the authors, we reused and adapted publicly available code: the supplementary material provided by Rodriguez and Laio (2014) [10] and the implementation from Gundersen et al. (2018) (available at https://github.com/AIReproducibility2018/ClusteringByFastSearchAndFindOfDensityPeaks). Once the copilot was satisfied, the paper and supplementary material was made available to the user we recruited for our reproducibility task. In practice, for the paper by Rodriguez and Laio (2014) [10], working with our system produced the data for each experiment and a Notebook containing functions to load each file in the right format for the paper’s method, the specific hyperparameter values for each of the experiments, and a short explanation on how to use the paper’s methods and analyze each result. These, together, were the missing pieces that made the reproduction of the paper take more than 30 hours in the study by Gundersen et al.
As a second feasibility test, we evaluated the coverage of the system’s content, which means its ability to identify and scaffold the core experimental content required for reproduction. To operationalize this, we focused on results presented in tables and figures, as these are clearer targets than smaller claims embedded in the manuscript text. This focus allowed us to systematically compare how much reproducible content the system is able to detect and guide users through.
To evaluate content coverage, we conducted a separate experiment in which we compared the Jupyter Notebook generated by our copilot with one that was manually created. This experiment focused on an Earth Sciences study, in part to test our prototype on another type of subject. Specifically, the study used Jupyter Notebooks [12] that were developed by Dr. Deborah Khider (one of the coauthors of this paper) to help teach the research behind a study by DeLong et al. (2011) [13]. The goal of these pedagogical Notebooks was not full reproducibility, but rather a guide for early-career scientists to learn to use data science libraries such as scikit-learn to perform a calibration task. While the original study focused on the robustness of the proxy signal in space and among different species of corals, the pedagogical Notebooks illustrate the use of Frequentist and Bayesian approaches in performing this calibration task. In this case, to mimic the original authors following the system’s suggestions, we adapted the code of Dr. Deborah Khider (available at https://khider.github.io/dry-tortugas-calibration-fun).
When comparing the Notebook produced by OpenPub with the pedagogical ones from Dr. Khider, we found that OpenPub provided a wide coverage of figures and tables, capturing not only the core figures and tables highlighted in the educational Notebooks, but also additional elements of the paper that are essential for full reproduction. As shown in Table 1, OpenPub covered all key figures and tables included in the educational Notebooks (green rows), as well as several additional visualizations and results that are necessary from a reproducibility point of view, but were excluded in the pedagogical version (white rows). The only omission from the coverage of OpenPub was a figure (Figure 8 in DeLong et al. (2011) [13]), which was also not prioritized in the teaching-oriented Notebooks (orange row). This comprehensive coverage demonstrates that such an AI system does not merely replicate what educators consider useful for instruction: it successfully identifies and scaffolds the full scope of reproducible content in a research paper, showing that this kind of tool can be quite complete when used for scientific validation.
Table 1: Comparative analysis of reproducibility coverage between OpenPub’s generated Notebook and Dr. Khider’s educational Notebooks [12] for the DeLong et al. (2011) study [13]. Green rows indicate figures and tables covered by both systems, orange rows mark a figure not covered by OpenPub but also excluded from the educational Notebooks, and white rows highlight additional figures and tables captured by OpenPub but omitted from the educational Notebooks.
Figures/Tables from DeLong et al. (2011) OpenPub’s Notebook Educational Notebooks
Table 2 $\surd$ $\times$
Figure 1 $\surd$ $\times$
Figure 2 $\surd$ $\surd$
Figure 3 $\surd$ $\times$
Figure 4 $\surd$ $\times$
Figure 5 $\surd$ $\surd$
Figure 6 $\surd$ $\times$
Figure 7 $\surd$ $\surd$
Figure 8 $\times$ $\times$
Figure 9 $\surd$ $\times$
Table 3 $\surd$ $\surd$
Discussion
During our research, we identified several challenges related to the use of AI to enhance reproducibility.
Scientific Understanding
Reproducibility can take many forms, such as computational reproducibility [14, 15, 16], empirical reproducibility [17], and outcome, analysis and interpretation reproducibility [18]. Throughout the use case in this paper, we focused on what can be called “rote reproducibility”, which can be defined as regenerating figures and tables using the same code and data. This type of reproducibility is related to a more complete “scientific reproducibility”, where the reconstruction of the reasoning and conclusions behind the research is sought. However, achieving complete scientific reproducibility is substantially more difficult as it requires access to the logic, rationale, and underlying assumptions that may not be explicitly stated in a paper.
In this case study, only preliminary steps were taken towards the challenging issues of scientific reproducibility. For example, we inferred analytical workflows from the manuscript and added preliminary instructional steps in the generated Notebook to guide readers in evaluating the plausibility of reproduced results. However, making it possible for the system to have a full understanding of the science behind the paper is a challenge to tackle in future work.
User Adaptation
A related complication comes from differences in the scientific background of the readers. What may be intuitive to an expert may be opaque to a student or newcomer. The tailoring of the guide based on the reader’s expertise is essential to promote scientific understanding. In future work, we plan to automatically construct semantic representations of the author analyses, enabling AI copilots to generate reader-specific guidance. This would not only support reproducibility, but also accelerate comprehension and learning for reviewers and readers alike. Indeed, our observations during our study suggest that simply understanding a paper often requires as much effort as reproducing the study using our Notebook.
Data Preparation and Diagnostic Analysis
Another prominent challenge is related to data formatting and preprocessing. When input data are not properly structured to work with the authors code, or worse, if essential preprocessing steps are missing, it becomes time-consuming for readers to identify and resolve these issues. At the same time, it remains difficult for an LLM to infer from raw numeric files that something is amiss. For instance, a file might appear valid on the surface, yet fail silently due to missing normalization. There are a number of ways to address this problem. For instance, if a paper explicitly states certain requirements on the input data (e.g., the data should have zero mean and unit variance), the system could test for these conditions and flag discrepancies. The system could also generate diagnostic code to validate whether the data meet the requirements specified by the authors. A related challenge is verifying that the output is reasonable/correct, i.e., analyzing the executed cells in the generated Notebook. While it may not always be feasible to verify whether generated results exactly match those in the original paper, at a minimum the system should at least detect major inconsistencies. For example, if the paper claims that five clusters were found in a dataset, but the system’s reproduction produces none, that discrepancy could prompt the system to recommend further investigation.
Claim Identification
A further challenge lies in identifying which claims in a manuscript warrant reproduction. While some are clearly marked in tables and figures - such as reported accuracy scores or performance benchmarks - others are embedded subtly in narrative text. Furthermore, different types of claim can coexist in the paper. Claims can be stated at the end of the paper and made based on the results of the study, or at the beginning to often be contrasted with each other and show where the new study adds knowledge. These latter types of claim often lead to the hypothesis to be tested in the paper. Finally, there are assumptions that are often made in scientific studies that could be presented as claims (e.g., “following the study by X et al., we will use the following method that has been proven valid for our data”). Building the capability to pinpoint all the claims, identify the ones that are reproducible, and verify that the supporting data and code are present is challenging.
Code Generation
Reproducing a study often (but not always) involves executing the code provided by the author. In some cases, there are bugs in that code, and in many cases, some parts of the code are missing, or at least, there is code that the reader needs to write to conveniently reproduce the work. Though in many cases we found that LLMs could do a fairly good job of automatically generating such code, we felt that this would be a distraction in this case study (especially if the generated code contained subtle errors), and we decided to rely on the authors to enter the code in the Notebook (who often actually has the code somewhere). In future work, we plan to reconsider how OpenPub can make use of the rapidly improving capabilities of modern LLMs to help with bug-free code generation as part of the reproducibility process.
Limitations of Evaluation
This study presents a feasibility study rather than a large-scale evaluation. Furthermore, our experiments assumed full compliance by the original authors with the system, which may not always be realistic in practice. Finally, the use of large language models introduces some variability due to non-deterministic outputs. These limitations underscore the preliminary nature of our findings, but also highlight clear directions for future work.
Conclusion
This use case has demonstrated the feasibility and potential of AI-driven copilot systems to meaningfully improve the openness of scientific research. We demonstrated the impact of such an AI-based approach, as well as its capacity to deliver actionable impact across different user roles, in particular authors and readers. As a concrete example, we focused this initial effort on reproducibility, an area of open science that remains difficult, yet critical to address.
As shown in the study by Gundersen et al. (2025), current barriers to reproducibility often stem from missing or unclear information [9]. In our evaluation, we showed that in a case that was documented in Gundersen et al., our AI system reduced the reproduction time from more than 30 hours to approximately 1 hour. This efficiency gain illustrates the value of AI-generated proactive guidance tailored to both authors and readers.
Importantly, the success of the system stems from its ability to surface specific, high-impact shortcomings - missing hyperparameters, inaccessible datasets, ambiguous code - and to communicate those deficiencies in ways that are actionable for authors. It also provides reproducibility scaffolds for readers, such as structured Jupyter Notebooks.
The broader implication of our findings is that reproducibility, often seen as a post-publication burden, can instead be integrated as a dynamic part of the scientific workflow. By helping authors anticipate reader needs and giving readers tools to navigate complex research materials, AI copilot systems can address complex problems by treating authors and readers as two faces of the same coin.
Although the system we developed for this use case focuses on “rote reproducibility” - the ability to regenerate results from the same datasets used in a study - future work should include pathways to support deeper, scientific reproducibility. These include modeling the logic and assumptions behind claims, adapting feedback to reader expertise, and even suggesting clarifications when results deviate.
Acknowledgments
This material is based upon work supported by the Office of Naval Research under Contract No. N68335-25-C-0016. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Office of Naval Research.